OpenStudy (anonymous):

f(x)=-5x+2,find f(-3)

OpenStudy (anonymous):

f(-3) = 17

OpenStudy (anonymous):

f(-3) = 17

OpenStudy (saifoo.khan):

f(-3) = 17

OpenStudy (anonymous):

f(x)=-5x+2,find f(-3) f(-3)=-5(-3)+2=15+2=17

OpenStudy (anonymous):

did u got it. sierra14